OpenStudy (babynini):

Find area of elipse using calculus. (4x^2)+y^2=9

OpenStudy (babynini):

What i've done so far \[f(x)=y=\sqrt{9-4x^2}\]\[x=(3/2)\sin(\theta)\]\[dx=(3/2)\cos(\theta)d \theta\]

OpenStudy (anonymous):

solve for y, then integrate in the upper right, then multiply by 4

OpenStudy (babynini):

\[4\int\limits_{0}^{1/2}\sqrt{9-4(\frac{ 3 }{ 2 }\sin(\theta))^2}*3\cos(\theta)d \theta\]

OpenStudy (babynini):

Yeah so I got that and then did a bunch of calculations and jazz and then at the end I got to \[36\int\limits_{0}^{1/2}\cos^2d \theta\]\[36(\frac{ \cos \theta \sin \theta+ \theta }{ 2 })|^{1/2}_0\] and that's whee i'm stuck. Because I don't want a decimal answer.

OpenStudy (babynini):

@satellite73

OpenStudy (anonymous):

wow that looks way complicated

OpenStudy (anonymous):

maybe we can do it an easier way, maybe not, but first off how did you get your limits of integration?

OpenStudy (babynini):

graphing xD

OpenStudy (babynini):

Well its not all that complicated because things cancel out pretty quickly..

OpenStudy (babynini):

oops!! I knew it was 1.5 I was just being dumb heh. 3/2 it is then xD

OpenStudy (anonymous):

maybe it is easier if we write it like an ellipse in the form of \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

in fact i know it is

OpenStudy (anonymous):

easiest of all is just to do it that way, get a formula with \(a,b\) in it

OpenStudy (babynini):

haha ok. but remember we have to use integrals and calc. I already did it using the area of an elipse and the prof said "must use integrals!"

OpenStudy (babynini):

yeaa using the pi*a*b would be epic. but not allowed :/

OpenStudy (anonymous):

yeah i mean derive the formula using integrals

OpenStudy (anonymous):

cause these constants are killing me

OpenStudy (babynini):

hmm. what do you think then? I mean, this way isn't too bad. Like everything cancels out. It's just the evaluating at 3/2 with sin and cos that I don't quite get.

OpenStudy (anonymous):

why not try \[f(x)=y=\sqrt{9-4x^2}=\sqrt{1-\frac{4}{9}x^2}\]

OpenStudy (anonymous):

then at least the integral will work nicer using \[x=\frac{2}{3}\sin(\theta)\]

OpenStudy (anonymous):

oh wait that is wrong, sorry

OpenStudy (anonymous):

i see the problem, hold on we can fix it

OpenStudy (babynini):

oh my gosh I see an error that I made :o

OpenStudy (babynini):

it's (3/2) cos not 3cos for the dx

OpenStudy (anonymous):

ok lets do it this way i will assume that all your work above is correct, since you already did it but you should change the limits of integration for sure, to 0 to \(\frac{\pi}{2}\)

OpenStudy (babynini):

ooo why can we change it to that?

OpenStudy (anonymous):

these constants are stressing me out, are you sure we can't use \(a\) and \(b\)?

OpenStudy (babynini):

got it. I got the answer, you genius you. Let me write out the process for you xD

OpenStudy (anonymous):

ok change the limits when you make the substitution

OpenStudy (anonymous):

btw i made a typo, meant to write \[f(x)=y=\sqrt{9-4x^2}=3\sqrt{1-\frac{4}{9}x^2}\]

OpenStudy (anonymous):

then make \(x=\frac{3}{2}\sin(\theta)\)

OpenStudy (anonymous):

then the change in limits is clear

OpenStudy (babynini):

\[y=\sqrt{9-4x^2}\]\[x=(3/2)\sin \theta\]\[dx=(3/2)\cos \theta d \theta\]\[4\int\limits_{0}^{\pi/2}\sqrt{9-4(\frac{ 3}{ 2 }\sin \theta)^2}*(3/2)\cos d \theta\]\[4\int\limits_{0}^{\pi/2} \sqrt{9-9\sin^2 \theta}*(3/2)\cos \theta d \theta\]\[4\int\limits_{0}^{\pi/2}3 \sqrt{1-\sin^2 \theta}*(3/2)\cos \theta d \theta\]\[18\int\limits_{0}^{\pi/2}\cos^2 \theta d \theta\]\[18(\frac{ \cos \theta \sin \theta + \theta }{ 2 })|^{\pi/2}_0\]\[18(\pi/2 - 0)\]\[\frac{ 9\pi }{ 2 }\]

OpenStudy (anonymous):

it is right?

OpenStudy (babynini):

Sorry, a bunch of little steps and calculations that I left out, but I trust you can follow that process. It takes a frikin amount of time to type those out heh Yeah that's the answer!

OpenStudy (babynini):

do I have to show how I got the limits of integration to be 0 and pi/2?

OpenStudy (anonymous):

i usually use \[\int\cos^2(\theta)d\theta =\frac{1}{2}\int(\cos(2\theta)+1)d\theta = \frac{1}{2}(\sin(2\theta)+\theta)\]

OpenStudy (anonymous):

no you don't i know how they got there

OpenStudy (anonymous):

the limits i mean you have \[x=\frac{3}{2}\sin(\theta)\] that gives the change in limits

OpenStudy (babynini):

ooh..should I use that then? I used this version of the answer for cos because I had done it in a previous question so I just wrote "see question 3" (we were asked to solve int of cos^2 by integration by parts)

OpenStudy (anonymous):

unless you are unclear about the limits?

OpenStudy (anonymous):

makes no difference i am sure

OpenStudy (babynini):

okay :) mm i'm a little unsure about the limits. But...do you think I would be expected to write out the process for finding the limit? It's a project homework type thing.

OpenStudy (anonymous):

lets do it, can't take but a second, and it will make it clear

OpenStudy (anonymous):

take \[x=\frac{3}{2}\sin(\theta)\] and solve for \(\theta \)

OpenStudy (anonymous):

i guess i should ask if you know how to do that

OpenStudy (babynini):

arcsin(2x/3) = theta ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if \(x=0\) you get arcsin(0)=0

OpenStudy (anonymous):

that is \[\theta=\sin^{-1}(0)=0\]

OpenStudy (anonymous):

and if \(x=\frac{3}{2}\) you get \[\theta =\sin^{-1}(\frac{2}{3}\times \frac{3}{2})=\sin^{-1}(1)=\frac{\pi}{2}\]

OpenStudy (anonymous):

now, if it was not clear before, you see where the \(\frac{\pi}{2}\) comes from right?

OpenStudy (babynini):

Yes! one little question just for clarification. We set x = 3/2 because that's the x intercept, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] it is \(\pm a\) in your case if you divide by 9 to put in in standard form you get the rather unwieldy \[\frac{x^2}{\left(\frac{3}{2}\right)^2}+\frac{y^2}{3^2}=1\]

OpenStudy (babynini):

oh I see :D

OpenStudy (anonymous):

ok good we done? all the i's dotted and t's crossed?

OpenStudy (babynini):

Absopositiveloutely. Thank you so much :)

OpenStudy (anonymous):

yw you did almost all the work!